In 2004 there were 6675 hybrid vehicles in use. In 2007, there were 7166 hybrid vehicles in use. Find the equation describing the relationship between time and number of hybrid vehicles. Make sure you use ordered pairs of the form (years past 1998, number of vehicles)
Y=

Answer by stanbon(48558)

In 2004 there were 6675 hybrid vehicles in use.
In 2007, there were 7166 hybrid vehicles in use.
---------------------------------------------------------
Find the equation describing the relationship between time and number of hybrid vehicles. Make sure you use ordered pairs of the form (years past 1998, number of vehicles)
----------------
Pairs: (6,6675) (9,7166)
slope = (7166-6675)/(9-6) = 163.67
intercept: 7166 = 163.67*9 + b; b = 5293
----------------------------
EQUATION:
Y= 163.67x + 5293
-------------------
Cheers,
Stan H.